# Resolving the Body-Order Paradox of Machine Learning Interatomic Potentials

### v1

@@ I Introduction
The many-body expansion (MBE) expresses the global observable of a chemical system as a sum of contributions from interactions at different body-orders, where the “bodies” are taken to be atoms, molecules, or larger fragments in the system. The MBE enables the interpretation of complex interactions in terms of simpler body-ordered contributions, and can provide reasonable approximations to the global quantity, especially when the contributions decay rapidly with the number of bodies involved. The MBE has given rise to many local or fragment-based quantum chemistry methods [1, 2, 3, 4, 5, 6, 7, 8, 9] that offer favorable scaling for large systems, as well as force fields that can be used for atomic scale simulations at large length and time scales [10, 11, 12].
Machine learning interatomic potentials (MLIPs) [13, 14, 15, 16, 17] enable ab initio-quality atomistic simulations with linear system size scaling and low prefactors, further extending the accessible length and time scales of the simulations. MLIPs commonly adopt a locality ansatz based on the nearsightedness principle [18, 19], and represent the local environments of chemical systems in terms of body-ordered correlations between the central atom and its neighbors. [20, 21, 22] This allows MLIPs to effectively capture the complex many-body interactions while retaining favorable scalability and transferability.
@@
In this study, we systematically analyze the body-orderedness of NN-based MLIPs and the resulting learning behavior. First, we revisit the trends in the MBE of ab initio calculations, which commonly serve as reference for MLIP training, at the atomic limit. Next, we train a curated set of MLIPs on datasets of hydrogen clusters to investigate how the models infer body-order trends of hydrogen [23, 24], and how they compare to the reference. We further explore how the trends vary between the different MLIPs and dataset makeup, and discuss how they may affect model accuracy and learning behavior. Finally, we consider the relationship between the body-ordered energetics and the out-of-distribution accuracy or generalizability of MLIPs.
@@ II.1 Parallelism between the MBE and the body-orderedness of MLIPs
$\varepsilon_{i}$ is the “local” energy associated with atom $i$, and it is a function of the local geometry described via $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$ for all $j$ in $A_{i}$, the local environment of atom $i$. The MLIPs differ in the functional form used to predict $\varepsilon_{i}$, which usually involves combining multiple $\mathbf{r}_{ij}$ terms to describe higher body-order correlations. [21, 25] Among the models, there exists a subset [15, 16] where features that formally depend on a fixed number of neighbors are used as the polynomial basis for a linear expansion:
@@
Here, $\nu$ is the “correlation order” involving multiple neighbor atoms $j$ (then, body order equals $\nu+1$). The expression tells us that the predicted energy can be explicitly decomposed into body-ordered contributions. Despite the apparent parallel with the MBE (Eq. (1)), previous works [26, 27, 28] have shown that there does not exist a 1-to-1 correspondence between the two because $\phi_{\nu}$ also contains terms associated with lower body-order correlations, unless they are explicitly elimintated [27] (see Appendix A).
In this study, we focus on the NN-based MLIPs, which have shown greater accuracies for much larger portions of the chemical space, [29, 30, 31, 32] yet their body-orderedness has never been considered in detail. For the NN-based models, $\varepsilon_{i}$ often involves nonlinearities that make it challenging to derive an analytical connection to the MBE. One of such models is the Behler-Parrinello NN (BPNN) [13], where the local descriptors of 2- and 3-body correlations are taken as inputs to a fully connected feed-forward network that predicts $\varepsilon_{i}$, which are then summed across the system to obtain the total energy. Here, we specifically consider SOAP-BPNN, which adopts the Smooth Overlap of Atomic Position (SOAP) [33] as the local descriptor of choice. Starting from the 3-body SOAP descriptors, SOAP-BPNN implicitly reaches higher body-orders through the multiple NN layers and the nonlinear activation functions. Since SOAP is an incomplete descriptor [34], the resulting SOAP-BPNN descriptor also exhibits incompleteness at higher body-orders.
More recent MLIPs incorporate message-passing NNs [35] and transformers [36] in their architectures. One example that we consider is MACE [37], an equivariant message-passing NN-based MLIP grounded in the atomic cluster expansion (ACE) formalism. [16] In MACE, each layer describes the body-order correlations of the given atomic environment up to $\nu=3$. Equivariant message passing effectively introduces a new one-particle basis, raising $\nu$ by 1. [25] With two of such layers, MACE explicitly achieves maximum body-order of 13 [38]. Lastly, we consider the point-edge transformer (PET) [39], a transformer-based model that does not enforce exact rotational symmetry. In PET, the softmax function used in the attention mechanism, as well as the activation functions in the feedforward blocks, implicitly lead to a theoretically infinite body-order.
@@ II.2 Hydrogen cluster sampling and calculation details
The 8-mers are sampled from the bulk hydrogen datasets of Cheng et al. [23] We construct two distinct datasets: a “high density” (high $\rho$) dataset in which the clusters are sampled from 100 configurations with the highest density (average $\rho=$ 1.34 g/cm3), and a “low density” (low $\rho$) dataset where the clusters are sampled from 100 configurations with the lowest density (average $\rho=$ 0.461 g/cm3). In the high $\rho$ dataset, each 8-mer is sampled by choosing a hydrogen at random from a given configuration, then taking its seven nearest neighbors. In the low $\rho$ dataset, sampling is performed by taking a random hydrogen atom and its closest neighbor, then adding three nearest pairs of atoms to complete the 8-mer. Example clusters are shown in Figure 1. Such difference in the sampling protocol ensures the sampling of distinct chemical trends seen in the corresponding bulk phases, where high $\rho$ systems are atomic, covalently bound and metallic, and low $\rho$ systems are molecular, bound by non-covalent interactions, and insulating. From each dataset, we randomly select 500 clusters, for which we also create the accompanying datasets containing canonically complete sets of their sub-clusters.
To obtain the energies and forces for MLIP training, DFT calculations are performed using FHI-aims [40]. The PBE exchange correlation functional [41] is employed with the “tight” species-default basis set, and Gaussian smearing with $\sigma=0.025$ eV is used to determine the occupations. To assess body-ordered energetics of hydrogen clusters at a much higher level of theory, we also perform spin-adapted density matrix renormalization group (DMRG) calculations [42] using Block2 [43] on a few representative clusters. Further details of the cluster sampling protocol and the reference calculations are given in the Supporting Information.
@@ II.3 MLIP training details
We consider three NN-based models widely varying in their architectures: SOAP-BPNN, MACE, and PET. In all cases, models are trained via stochastic gradient descent with the Adam optimizer [44] on the mean squared loss of energies and forces. Train, validation, and test splits of the datasets are kept consistent between the models. We set a consistent cutoff radius of 5.5 Å in all three models, which corresponds to the value used in the MLIPs developed in the previous work and encompasses the entire cluster. [23] In SOAP-BPNN and MACE, the baseline energy, i.e. the energy of the isolated atom as perceived by the model, is fixed to the reference isolated atom energy. In PET, the existence of central tokens and their incorporation into the attention mechanism prevents the model from having a strict, pre-defined baseline. Instead, we take a data-driven approach and include an isolated atom reference configuration in the training sets for PET, allowing the model to learn an isolated atom energy to a good accuracy (<0.01<0.01 eV from the reference value). We note that any deviation in the learned isolated atom energy from the reference may give PET a small advantage in the learning tasks. The ratio between the energy and force losses is fixed at 1:1 for SOAP-BPNN and PET. For MACE, we adopt its default training routine, which is to use 1:100 ratio between energy and force losses for the first stage, and 1000:100 in the second stage. The rest of the training details and hyperparameters are set to the defaults of the models. Full sets of the inputs and hyperparameters are provided in our Materials Cloud repository. [45]
@@ III Body-ordered energetics of hydrogen clusters
Before any MLIP analysis, we first evaluate the body-ordered energetics of hydrogen 8-mers in the reference ab initio calculations. In many cases, the MBE formalism is applied at the molecular level [46, 47, 48], where $V^{(m)}$ is expected to converge quickly to zero with increasing $m$. An archetypal system is water [49, 50, 51, 52, 53, 54, 55, 56, 57, 58], for which many-body force fields have shown good accuracies with limited body-orders, up to 4, including also a polarizable model baseline. Even for water, however, body-ordered energetics at the DFT theory-level can converge slowly, and show large oscillations when ions are introduced. [59, 60] At the atomic level, where covalent and metallic interactions dominate, even more complex and non-trivial trends have been observed for mercury, sodium, silicon, and gold. [61, 62]
@@
In both cases, an oscillatory and divergent trend is observed in the DFT body-ordered energetics (black markers). A negative energy contribution is first observed at $m=2$, then a larger positive contribution at $m=3$, and the sign of $\tilde{V}^{(m)}_{A}$ continues to alternate with the magnitude increasing with $m$. The anticipated chemical trend between the low $\rho$ and high $\rho$ 8-mers is manifested as a difference in the magnitudes of $\tilde{V}^{(m)}_{A}$ across all $m$, with larger contributions observed for the high $\rho$ 8-mers over the low $\rho$ 8-mers. Similar trends persist at the full configuration interaction (FCI) theory level using DMRG, revealing that the apparent trend is not a mere consequence of the approximate nature of DFT [60] (see Appendix B for further consideration of the DFT self-interaction error).
Previously, long-range many-body interactions have been observed in 1D hydrogen chains of up to 50 atoms [63], where strong antiferromagnetic (AFM) spin-spin correlations are present. Here, we also analyze spin-spin correlations $C_{ij}$ for representative 8-mers in three spatial dimensions, where $C_{ij}=\langle\hat{n}_{i\uparrow}\hat{n}_{j\downarrow}\rangle-\langle\hat{n}_{i\uparrow}\rangle\langle\hat{n}_{j\downarrow}\rangle$. Figure 1 reveals that both low and high $\rho$ 8-mers display substantial spin-spin correlations across all atomic pairs, thereby contributing to the apparent non-convergence of the body-ordered energetics. The correlation patterns reveal distinct behaviors between the two density regimes: in the low $\rho$ 8-mer, strong correlations are present between bonded atom pairs, and non-negligible correlations also persist between the intermolecular pairs. The high $\rho$ 8-mer exhibits stronger and more delocalized spin correlations across all pairs within the cluster.
These results demonstrate that the “true” body-ordered energetics of hydrogen 8-mers are inherently oscillatory and non-converging. We note that the contrast between our results and the mathematically proven exponential convergence of body-orders [64] arises from the choice of $E_{1}$ in the MBE. While our expansion references itself to the “vacuum” (i.e., each sub-cluster is considered in complete isolation), considerations in Ref. [64] assume full awareness of the entire local environment even for the sub-clusters. In Appendix C, we further unravel the dependence of the convergence trends on the baseline choice, and rationalize that for energy-stable atomic clusters such as the hydrogen 8-mers of our study, the observed oscillatory trend in the vacuum-referenced MBE is reasonable under the choice of $E_{1}$ as the isolated atom energy.
@@ IV Effective body-orderedness of MLIPs
Figure 2 reveals that all MLIPs deviate from the DFT body-order trends in both energies and forces and infer their own, effective body-orders for the hydrogen 8-mers, with much smaller magnitudes across all $m$. In SOAP-BPNN, body-orders of the low $\rho$ 8-mers are fast-converging with $m$, whereas those of the high $\rho$ 8-mers exhibit an oscillatory, slow-converging behavior in $\tilde{V}^{(m)}_{A}$ and an overall increasing trend in $|\partial\tilde{V}^{(m)}_{A}/\partial\textbf{r}|$. The spread of values across individual samples is also much narrower for low $\rho$ 8-mers compared to the high $\rho$ 8-mers that reach a standard deviation ($\sigma$) of 0.481 eV and 3.704 eV/Å for $m=8$.
@@
In Figure S1, we present the “per-mm” RMSEs of the three MLIPs on the sub-clusters of the test set 8-mers, assessing the accuracy of the effective body-ordered energetics of the MLIPs. All three models exhibit similarly large RMSEs on average for both energies (325 meV/atom) and forces (1.21 eV/Å) for all $m<8$. This corroborates that the intuited body-ordered energetics are only effective, and significantly far from the DFT reference.

### v2

@@ I Introduction
The many-body expansion (MBE) expresses the global observable of a chemical system as a sum of contributions from interactions at different body-orders, where the “bodies” are taken to be atoms, molecules, or larger fragments in the system. The MBE enables the interpretation of complex interactions in terms of simpler body-ordered contributions, and can provide reasonable approximations to the global quantity, especially when the contributions decay rapidly with the number of bodies involved. The MBE has given rise to many local or fragment-based quantum chemistry methods [Yang1991, Saebo1993, Yang1995, Baer1997, Kitaura1999, Dahlke2007, Solovyev2008, Richard2012, Pruitt2014] that offer favorable scaling for large systems, as well as force fields that can be used for atomic scale simulations at large length and time scales [Rappe1992, Wang2004, Vanomm2010].
Machine learning interatomic potentials (MLIPs) [Behler2007, Bartok2010, Shapeev2016, drau19prb, Behler2021] enable ab initio-quality atomistic simulations with linear system size scaling and low prefactors, further extending the accessible length and time scales of the simulations. MLIPs commonly adopt a locality ansatz based on the nearsightedness principle [Kohn1996, Prodan2005], and represent the local environments of chemical systems in terms of body-ordered correlations between the central atom and its neighbors. [Willatt2018, musi+21cr, Glielmo2021] This allows MLIPs to effectively capture the complex many-body interactions while retaining favorable scalability and transferability.
@@
In this study, we systematically analyze the body-orderedness of NN-based MLIPs and the resulting learning behavior. First, we revisit the trends in the MBE of ab initio calculations, which commonly serve as reference for MLIP training, at the atomic limit. Next, we train a curated set of MLIPs on datasets of hydrogen clusters to investigate how the models infer body-order trends of hydrogen [Cheng2020, Bischoff2024], and how they compare to the reference. We further explore how the trends vary between the different MLIPs and dataset makeup, and discuss how they may affect model accuracy and learning behavior. Finally, we consider the relationship between the body-ordered energetics and the out-of-distribution accuracy or generalizability of MLIPs.
@@ II.1 Parallelism between the MBE and the body-orderedness of MLIPs
$\varepsilon_{i}$ is the “local” energy associated with atom $i$, and it is a function of the local geometry described via $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$ for all $j$ in $A_{i}$, the local environment of atom $i$. The MLIPs differ in the functional form used to predict $\varepsilon_{i}$, which usually involves combining multiple $\mathbf{r}_{ij}$ terms to describe higher body-order correlations. [musi+21cr, Nigam2022] Among the models, there exists a subset [Shapeev2016, drau19prb] where features that formally depend on a fixed number of neighbors are used as the polynomial basis for a linear expansion:
@@
Here, $\nu$ is the “correlation order” involving multiple neighbor atoms $j$ (then, body order equals $\nu+1$). The expression tells us that the predicted energy can be explicitly decomposed into body-ordered contributions. Despite the apparent parallel with the MBE (Eq. (1)), previous works [Chong2025, Ho2024, Tan2025] have shown that there does not exist a 1-to-1 correspondence between the two because $\phi_{\nu}$ also contains terms associated with lower body-order correlations, unless they are explicitly elimintated [Ho2024] (see Appendix A).
In this study, we focus on the NN-based MLIPs, which have shown greater accuracies for much larger portions of the chemical space, [Batatia2024, Park2024, Mazitov2025, Rhodes2025] yet their body-orderedness has never been considered in detail. For the NN-based models, $\varepsilon_{i}$ often involves nonlinearities that make it challenging to derive an analytical connection to the MBE. One of such models is the Behler-Parrinello NN (BPNN) [Behler2007], where the local descriptors of 2- and 3-body correlations are taken as inputs to a fully connected feed-forward network that predicts $\varepsilon_{i}$, which are then summed across the system to obtain the total energy. Here, we specifically consider SOAP-BPNN, which adopts the Smooth Overlap of Atomic Position (SOAP) [Bartok2013] as the local descriptor of choice. Starting from the 3-body SOAP descriptors, SOAP-BPNN implicitly reaches higher body-orders through the multiple NN layers and the nonlinear activation functions. Since SOAP is an incomplete descriptor [Pozdn2020], the resulting SOAP-BPNN descriptor also exhibits incompleteness at higher body-orders.
More recent MLIPs incorporate message-passing NNs [Gilmer2017] and transformers [Vaswani2023] in their architectures. One example that we consider is MACE [bata+22nips], an equivariant message-passing NN-based MLIP grounded in the atomic cluster expansion (ACE) formalism. [drau19prb] In MACE, each layer describes the body-order correlations of the given atomic environment up to $\nu=3$. Equivariant message passing effectively introduces a new one-particle basis, raising $\nu$ by 1. [Nigam2022] With two of such layers, MACE explicitly achieves maximum body-order of 13 [Kovacs2023]. Lastly, we consider the point-edge transformer (PET) [Pozdn2023], a transformer-based model that does not enforce exact rotational symmetry. In PET, the softmax function used in the attention mechanism, as well as the activation functions in the feedforward blocks, implicitly lead to a theoretically infinite body-order.
@@ II.2 Hydrogen cluster sampling and calculation details
The 8-mers are sampled from the bulk hydrogen datasets of Cheng et al. [Cheng2020] We construct two distinct datasets: a “high density” (high $\rho$) dataset in which the clusters are sampled from 100 configurations with the highest density (average $\rho=$ 1.34 g/cm3), and a “low density” (low $\rho$) dataset where the clusters are sampled from 100 configurations with the lowest density (average $\rho=$ 0.461 g/cm3). In the high $\rho$ dataset, each 8-mer is sampled by choosing a hydrogen at random from a given configuration, then taking its seven nearest neighbors. In the low $\rho$ dataset, sampling is performed by taking a random hydrogen atom and its closest neighbor, then adding three nearest pairs of atoms to complete the 8-mer. Example clusters are shown in Figure 1. Such difference in the sampling protocol ensures the sampling of distinct chemical trends seen in the corresponding bulk phases, where high $\rho$ systems are atomic, covalently bound and metallic, and low $\rho$ systems are molecular, bound by non-covalent interactions, and insulating. From each dataset, we randomly select 500 clusters, for which we also create the accompanying datasets containing canonically complete sets of their sub-clusters.
To obtain the energies and forces for MLIP training, DFT calculations are performed using FHI-aims [Blum2009]. The PBE exchange correlation functional [PBE] is employed with the “tight” species-default basis set, and Gaussian smearing with $\sigma=0.025$ eV is used to determine the occupations. To assess body-ordered energetics of hydrogen clusters at a much higher level of theory, we also perform spin-adapted density matrix renormalization group (DMRG) calculations [Sharma2012] using Block2 [zhai2023block2] on a few representative clusters, using one- and two-electron integrals generated by PySCF [sun2018pyscf, sun2020recent]. During method investigation, coupled-cluster with single, double, and triple excitations (CCSDT) calculations were also carried out in Q-Chem [epifanovsky2021software] for internal checks. Further details of the cluster sampling protocol and the reference calculations are given in the Supplementary Material.
@@ II.3 MLIP training details
We consider three NN-based models widely varying in their architectures: SOAP-BPNN, MACE, and PET. In all cases, models are trained via stochastic gradient descent with the Adam optimizer [Kingma2017] on the mean squared loss of energies and forces. Train, validation, and test splits of the datasets are kept consistent between the models. We set a consistent cutoff radius of 5.5 Å in all three models, which corresponds to the value used in the MLIPs developed in the previous work and encompasses the entire cluster. [Cheng2020] In SOAP-BPNN and MACE, the baseline energy, i.e. the energy of the isolated atom as perceived by the model, is fixed to the reference isolated atom energy. In PET, the existence of central tokens and their incorporation into the attention mechanism prevents the model from having a strict, pre-defined baseline. Instead, we take a data-driven approach and include an isolated atom reference configuration in the training sets for PET, allowing the model to learn an isolated atom energy to a good accuracy (<0.01<0.01 eV from the reference value). We note that any deviation in the learned isolated atom energy from the reference may give PET a small advantage in the learning tasks. We have seen, however, that allowing the MLIPs to adjust freely their baseline energy does not impact significantly our observations (see Supplementary Material). The ratio between the energy and force losses is fixed at 1:1 for SOAP-BPNN and PET. For MACE, we adopt its default training routine, which is to use 1:100 ratio between energy and force losses for the first stage, and 1000:100 in the second stage. The rest of the training details and hyperparameters are set to the defaults of the models. Full sets of the inputs and hyperparameters are provided in our Materials Cloud repository. [Talirz2020]
@@ III Body-ordered energetics of hydrogen clusters
Before any MLIP analysis, we first evaluate the body-ordered energetics of hydrogen 8-mers in the reference ab initio calculations. In many cases, the MBE formalism is applied at the molecular level [Riera2020, Riera2020b, Bull2023], where $V^{(m)}$ is expected to converge quickly to zero with increasing $m$. An archetypal system is water [Gora2011, Babin2012, Medders2013, Medders2013b, Richard2014, Pinski2014, Lao2016, Paesani2016, Heindel2021, Zhai2023], for which many-body force fields have shown good accuracies with limited body-orders, up to 4, including also a polarizable model baseline. Even for water, however, body-ordered energetics at the DFT theory-level can converge slowly, and show large oscillations when ions are introduced. [Ouyang2014, Broderick2024] In the context of MLIPs and their body-orders, where body-ordered correlations or other geometric features are computed between atoms, the consistent scale at which to apply the MBE formalism is also that of individual atoms. At the atomic level, where covalent and metallic interactions dominate, even more complex and non-trivial trends have been previously observed for mercury, sodium, silicon, and gold. [Paulus2004, Hermann2007]
@@
In both cases, an oscillatory and divergent trend is observed in the DFT body-ordered energetics (black markers). A negative energy contribution is first observed at $m=2$, then a larger positive contribution at $m=3$, and the sign of $\tilde{V}^{(m)}_{A}$ continues to alternate with the magnitude increasing with $m$. The anticipated chemical trend between the low $\rho$ and high $\rho$ 8-mers is manifested as a difference in the magnitudes of $\tilde{V}^{(m)}_{A}$ across all $m$, with larger contributions observed for the high $\rho$ 8-mers over the low $\rho$ 8-mers. Similar trends persist with the exact correlation treatment using DMRG, revealing that the apparent trend is not a mere consequence of the approximate nature of DFT [Broderick2024] (see Appendix B for further consideration of the DFT self-interaction error).
Previously, long-range many-body interactions have been observed in 1D hydrogen chains of up to 50 atoms [Motta2020], where strong antiferromagnetic (AFM) spin-spin correlations are present. Here, we also analyze spin-spin correlations $C_{ij}$ for representative 8-mers in three spatial dimensions, where $C_{ij}=\langle\hat{n}_{i\uparrow}\hat{n}_{j\downarrow}\rangle-\langle\hat{n}_{i\uparrow}\rangle\langle\hat{n}_{j\downarrow}\rangle$. Figure 1 reveals that both low and high $\rho$ 8-mers display substantial spin-spin correlations across all atomic pairs, thereby contributing to the apparent non-convergence of the body-ordered energetics. The correlation patterns reveal distinct behaviors between the two density regimes: in the low $\rho$ 8-mer, strong correlations are present between bonded atom pairs, and non-negligible correlations also persist between the intermolecular pairs. The high $\rho$ 8-mer exhibits stronger and more delocalized spin correlations across all pairs within the cluster.
These results demonstrate that the “true” body-ordered energetics of hydrogen 8-mers are inherently oscillatory and non-converging. We note that the contrast between our results and the mathematically proven exponential convergence of body-orders [Thomas2022] arises from the choice of $E_{1}$ in the MBE. While our expansion references itself to the “vacuum” (i.e., each sub-cluster is considered in complete isolation), considerations in Ref. [Thomas2022] assume full awareness of the entire local environment even for the sub-clusters, which is more consistent with the stated goal of investigating the convergence of MLIPs, even though it leaves the definition of $E_{1}$ somewhat vague. In Appendix C, we further unravel the dependence of the convergence trends on the baseline choice, and rationalize that for energy-stable atomic clusters such as the hydrogen 8-mers of our study, the observed oscillatory trend in the vacuum-referenced MBE is reasonable under the choice of $E_{1}$ as the isolated atom energy.
@@ IV Effective body-orderedness of MLIPs
Figure 2 reveals that all MLIPs deviate from the DFT body-order trends in both energies and forces and infer their own, effective body-orders for the hydrogen 8-mers, with much smaller magnitudes across all $m$. In SOAP-BPNN, body-orders of the low $\rho$ 8-mers are fast-converging with $m$, whereas those of the high $\rho$ 8-mers exhibit an oscillatory, slow-converging behavior in $\tilde{V}^{(m)}_{A}$ and an overall increasing trend in $|\partial\tilde{V}^{(m)}_{A}/\partial\textbf{r}|$. The spread of values across individual samples is also much narrower for low $\rho$ 8-mers compared to the high $\rho$ 8-mers that reach a standard deviation ($\sigma$) of 0.481 eV and 3.704 eV/Å in energies and forces for $m=8$.
@@
In Figure S1, we present the “per-mm” RMSEs of the three MLIPs on the sub-clusters of the test set 8-mers, assessing the accuracy of the effective body-ordered energetics of the MLIPs. All three models exhibit similarly large RMSEs on average for both energies (0.325 eV/atom) and forces (1.21 eV/Å) for all $m<8$. This corroborates that the intuited body-ordered energetics are only effective, and significantly far from the DFT reference. In the Supplementary Material, we also explore the body-order trends of the MLIPs when $E_{1}$ is no longer fixed to the isolated atom energy. We learn that all MLIPs retain similar trends with notably smaller contributions across all body-orders, with the exception of PET exhibiting slightly larger $|\partial\tilde{V}^{(m)}_{A}/\partial\textbf{r}|$ values for both low $\rho$ and high $\rho$ 8-mers.